OpenStudy (anonymous):

Given: F(x)=xsq-7 and given f(x)=|x|, g(x)=x-7, and h(x)=x sq it is asked to write each of the following functions as in F(x)= x sq-7 as a composition of the functions chosen from f,g,h, I guess this is kind of working backwards but I don't have a clue to start. Please show me your work. Thank you.

OpenStudy (anonymous):

By the sq do you mean square root? Something else? Also, from what you've posted, it doesn't seem entirely clear what you need to do. I assume you need to find the function composition using f,g, and h to get F?

OpenStudy (anonymous):

This is my first time out with this type of site. How do I acess a reply?

OpenStudy (anonymous):

oh my actually sq should be as in x squared

OpenStudy (anonymous):

Aylin am I still with you? Is there an answer and if so how do I find it? thank you

OpenStudy (anonymous):

Ok, so you have\[f(x)=|x|\]\[g(x)=x-7\]\[h(x)=x^{2}\]and you need to get\[F(x)=x^{2}-7\]out of it. Function composition is of the following form:\[(a^{o}b)(x)=a(b(x))\] Essentially, what you are doing with the function a is putting in b(x) wherever x is. So if a(x)=7x+4, and b(x)=x-2, then a(b(x))=7(b(x))+4=7(x-2)+4=7x-10. So what you need first is a function that turns x into x^2. Then you need a function that turns x into x-7. Does this make sense?

OpenStudy (anonymous):

Not yet Aylin can you give me a bit more?

OpenStudy (anonymous):

Hmm, perhaps it would help if you went through a couple of the function compositions... Can you find f(g(x)) and h(g(x)) for example?

OpenStudy (anonymous):

not sure I can. I mean f would be the answer if g(x) would be the value then that value would be applied to f(x)

OpenStudy (anonymous):

\[f(g(x))=|g(x)|=|x-7|\]\[h(g(x))=(g(x))^{2}=(x-7)^{2}=x^{2}-14x+49\]Does this make more sense now?

OpenStudy (anonymous):

so if I start with x sq -14x +49 am I suposed to be able to derive this back to h(g(x)) Given the f(x)=|x| and the other values of g and h?

OpenStudy (anonymous):

If you factor\[x^{2}-14x+49\]you get\[(x-7)\times (x-7)=(x-7)^{2}\]Then you know that you need a function that given an input will give an output of 7 less than the input, and another function that squares the input. So if you have g(x)=x-7 and h(x)=x^2, you can see that h(g(x))=(g(x))^2=x^2-14x+49, so that leads you back to h(g(x)).

OpenStudy (anonymous):

This makes sense to factor . I need to try this on some other problems and see what comes up also I just found the ^ sign for square. This is new to me ie online help so I must thank you soooo much. I will more than likely be back but I don't want to tie you up for now Berthacow

OpenStudy (anonymous):

I'm doing other things while I help you. (Actually I'm just watching TV).

OpenStudy (anonymous):

I am doing this on my own no school an it takes my full concentration wish it were more clear so thank you again.

OpenStudy (anonymous):

You're welcome. It's no problem at helping. Does the process make sense now though? If it helps you can ignore f(x) entirely for the problem you posted.

OpenStudy (anonymous):

(This is because h(f(x))=|x|^2=x^2, which is just h(x).)

OpenStudy (anonymous):

Alternatively, f(h(x))=|x^2|=x^2 since x^2 is necessarily positive for real numbers.

OpenStudy (anonymous):

What I figure to do is to print this all up and so to speak use it as a template to see if it works on the other problems and see what happens.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok you've been very kind go back to your TV and Thank you Berthacow out